# What Makes Good Data for Alignment? A Comprehensive Study of Automatic Data Selection in Instruction Tuning

### v1

@@ 1 Introduction
Figure 1: Illustration of the data selection approach. We measure data from three dimensions: complexity, quality, and diversity. $I$ and $R$ represent instruction and response respectively. For Evol Complexity and Evol Quality, we first collect samples with varying complexities or qualities through adopting an evolution-based approach following Xu et al. (2023), then we ask ChatGPT to rank and score the variants of the same data sample for a small seed dataset, and we train our own complexity and quality scorers based on these scores. In the last step, we utilize the trained scorers and adopt a score-first, diversity-aware approach to select the “good” data samples, as we detail in §3.
In the development of large language models (LLMs), aligning LLMs with human preferences is a necessary step for the models to accurately understand human instructions and generate relevant responses. Standard approaches for LLM alignment encompass instruction tuning and reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022). Instruction tuning, or supervised fine-tuning (SFT), refines the pre-trained model using annotated instructional data, often serving as the foundational step before RLHF to facilitate the model’s initial alignment (Ouyang et al., 2022; Touvron et al., 2023b). RLHF, on the other hand, leverages reinforcement learning to train models based on annotated feedback on their generated responses. While RLHF has underpinned the development of ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023), recent researches suggest that instruction tuning in isolation can offer competitive results (Sun et al., 2023; Zhou et al., 2023).
@@
In this work, we seek to define the characteristics of “good data” for instruction tuning, based on which we aim to push data efficiency of instruction tuning further in an automatic manner. To this end, we first explore various methods to quantitatively assess data examples from three key dimensions: complexity, quality, and diversity. Our hypothesis posits that the most effective datasets for instruction tuning are those that are complex, of high quality, and diverse. We examine a wide range of data measurement baselines and propose new metrics, which show a stronger correlation with alignment performance after instruction tuning. Drawing inspiration from Xu et al. (2023), we introduce Evol Complexity and Evol Quality, which evolve a single data point to produce a series of examples varying in complexity or quality, then we rand and score a small subset of them with ChatGPT and train a scorer to predict these scores, as illustrated in Figure 1. Such an evolution-based method automatically yields examples sorted based on complexity or quality, enabling enhanced and finer-grained data measurements. Combined with our diversity metric via distance of model embeddings, we design a simple strategy to select the most effective data examples from a large data pool. Being simple and effective, our data selection strategy paves the way for a paradigm of data-efficient instruction tuning – where fewer training samples can be automatically selected out to yield performance on par with, or even surpassing, models trained on significantly larger datasets.
We present Deita (short for Data-Efficient Instruction Tuning for Alignment), a family of models fine-tuned from the LLaMA (Touvron et al., 2023a; b) and Mistral (Jiang et al., 2023) models using our proposed technique to maximize data efficiency. On experiments in terms of MT-bench (Zheng et al., 2023), AlpacaEval (Li et al., 2023c), and the Open LLM Leaderboard (Beeching et al., 2023) demonstrate that Deita is able to outperform or be on par with state-of-the-art instruction following models such as Zephyr (Tunstall et al., 2023), Vicuna (Chiang et al., 2023) and WizardLM (Xu et al., 2023), while using over 10x fewer automatically selected data examples. For example, Deita-Mistral-7B – our model based on Mistral-7B – achieves 7.22 MT-bench and 80.78% AlpacaEval scores when trained on only 6K data samples with vanilla SFT training. After equipped with direct preference optimization (DPO,  Rafailov et al. (2023)), Deita-Mistral-7B + DPO, trained with 6K SFT and 10K DPO samples, obtains 7.55 MT-Bench and 90.06% AlpacaEval scores. In addition to the Deita model checkpoints, we release our light yet effective SFT datasets to facilitate efficient alignment for future researches.
@@ 2.2 Experimental Setup
To investigate data selection from large data pools, we construct two data pools with distinct properties to mimic different practical settings: (1) $X_{sota}$, which is constructed by ensembling the training datasets of the state-of-the-art aligned LLMs. This represents the setting where a data pool that is relatively complex, diverse, and of high-quality is available, and we aim to further improve the data efficiency in this case. Specifically, we follow Lu et al. (2023) to combine the datasets WizardLM (Alpaca), WizardLM (ShareGPT) (Xu et al., 2023), UltraChat (Ding et al., 2023), and ShareGPT (Chiang et al., 2023), resulting in a dataset of  300K samples; and (2) $X_{base}$, which mimics the scenario where the available data pool is overall lower-quality and redundant. This may be closer to the actual setting that people would encounter in practice. For $X_{base}$, we have utilized data from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), OAssit (Köpf et al., 2023), and FLAN 2022 (Longpre et al., 2023) to construct a dataset of 100K samples. Such quality-based separation of these open-source instruction tuning datasets roughly aligns with the analysis results in Lu et al. (2023); Li et al. (2023b). Statistics of the two data pools are summarized in Table 1.
@@
In the empirical study of this section, we fine-tune the LLaMA-1 13B model (Touvron et al., 2023b) on the instruction tuning dataset unless otherwise specified. We assume a data budget $m$ of 6K examples. We use the same hyperparameters for all the training as detailed in Appendix A. To evaluate alignment performance, we utilize MT-Bench (Zheng et al., 2023), a challenging benchmark that is commonly adopted to assess the instruction-following ability of LLMs. MT-bench consists of multi-turn conversations across various domains such as writing, reasoning, math, and coding. GPT-4 is employed as the judge to score model responses in MT-bench, which was found to produce a high agreement with humans (Zheng et al., 2023).
@@
Table 2: MT-bench scores of different complexity measurements. All methods select 6K samples. “Pool=50K” denotes the data selection procedure is conducted in a 50K-sized subset due to the cost of using ChatGPT to annotate the entire pool. We include the results of our method on the 50K data pool as well to make a fair comparison with the respective baselines.
@@ 2.3 From the Complexity Perspective – Evol Complexity
It is often believed that long, difficult, and complex data samples are more beneficial for instruction tuning (Zhao et al., 2023; Cao et al., 2023). For example, Xu et al. (2023) prompt ChatGPT to “evolve” data samples to deliberately increase their complexities, which leads to the state-of-the-art open-source alignment models WizardLM. In this section, we systematically study various metrics to assess the complexity of data samples and aim to identify the notion of complexity that contributes the most to the instruction-following ability of models. Concretely, we only consider the complexity dimension and define the selection strategy $\pi_{complexity}$ as selecting the $m$ examples with the highest complexity scores.
@@
As illustrated in the left part of Figure  1, we then ask ChatGPT to rank and score these 6 samples (prompt in Appendix E.2), obtaining the complexity scores $c$ corresponding to the instructions. We emphasize that, distinct from direct scoring, we give ChatGPT all 6 samples within one prompt – these samples represent different evolution stages of the same original sample and such a scoring scheme helps ChatGPT capture the small complexity differences among them, which leads to complexity scores to achieve finer-grained complexity differentiation among samples. We find that this is critical since otherwise, ChatGPT tends to assign similar scores to most examples, as we showcase in Appendix B. After obtaining ChatGPT scores on the small seed dataset, we use the scores to train an LLaMA-1 7B model to predict the complexity score given the input instruction. In the case of multi-turn dialogues, we score each turn separately and use the sum of them as the final score. In the experiments across this paper, we use 2K examples randomly sampled from the Alpaca dataset (Taori et al., 2023) as the seed dataset.
@@ Results:
Table 2 presents the results of selecting the 6K data samples from $X_{sota}$ and $X_{base}$ using various complexity metrics. Our Evol-Complexity metric leads to the best alignment performance on MT-bench. We observe that while Instag Complexity performs well on $X_{sota}$, its performance on $X_{base}$ is only slightly better than Random Selection. In contrast, Evol Complexity achieves superior performance on both datasets, indicating strong robustness across different dataset pools. Additionally, our method has significant advantages compared to methods that heavily rely on annotating with ChatGPT, such as Direct Scoring and Instruction Node. The experimental results also suggest that instruction lengths are not good indicators of preferred data by alignment. Interestingly, perplexity, an intuitive measure for complexity, produces much worse results than the random selection baseline. Through further investigation, we found that samples with large perplexity typically exhibit very short responses.
@@ 2.4 From the Quality Perspective – Evol Quality
In a manner akin to Evol Complexity, we introduce Evol Quality to augment the discernment of quality measurement. For a given data sample $(I_{k}^{(0)},R_{k}^{(0)})$, we prompt ChatGPT to elevate the quality of the response in an evolved way (please refer to Appendix E.4 for the prompt). This primarily involves enhancing helpfulness, augmenting relevance, enriching depth, fostering creativity, and supplying additional details. After $M$ iterations, for the same instruction $I_{k}^{(0)}$, we procure a set of responses spanning various qualities for $R_{k}$, denoted as $\{R_{k}^{(0)},\cdots,R_{k}^{(M)}\}$. Similar to Evol Complexity we set $M$ to 5.
@@ Results:
Table 3: MT-bench scores of different quality measurements. All methods select 6K samples for training. “Pool=50K” denotes the data selection procedure is conducted in a 50K-sized subset due to the cost of using ChatGPT to annotate the entire pool. We include the results of our method on the 50K data pool as well to make a fair comparison with the respective baselines.
@@ 2.5 From the Diversity Perspective – An Embedding-based Approach
We examine our embedding-based approach, which we refer to as Repr Filter. Specifically, we take the distance between sample $x_{i}$ and its nearest neighbor in $S$ as the metric to define $\mathcal{F}$. We leverage the LLaMA-1 13B model to encode the sentence and compute the cosine distance $d$, then $\mathcal{F}\coloneqq d>\tau$, where $\tau\in(0,1)$ is a threshold hyperparameter. This means that we consider an example $x_{i}$ could increase the diversity of $S$ when the embedding distance between $x_{i}$ and its nearest neighbor is larger than a threshold. During data selection, we first sort the data pool $X$ according to complexity and quality scores that we will detail in §3.1, then we examine each sample one by one and put $x_{i}$ to $S$ if $d>\tau$, where $S$ is initialized as empty. This process stops until the size of $S$ reaches the data budget $m$. We set threshold $\tau$ as $0.9$ in all the relevant experiments across this paper, while we provide analysis results on $\tau$ and different sentence representations in Appendix C.1.
@@ 3.1 Method
We train Deita by fine-tuning the pretrained models using the selected dataset with $m$ samples. We denote the resulting models as $\textsc{Deita}_{m}$. In this paper, we train Deita models based on LLaMA-1-13B, LLaMA-2-13B, and Mistral-7B (Jiang et al., 2023) respectively, and the training details are described in Appendix A.
@@
[…]
Table 5: Comparison of different data selection approaches, the backbone model is LLaMA-1-13B. For Alpagasus, we could not use ChatGPT to score all the examples in the pool due to the cost, thus we only score 50K random examples and select. $\dagger$ denotes the results obtained by using their released LLaMA-7B tagger model for a fair comparison.
@@
[…]
3.2 Experimental Setup
We train the Deita models with data budget of 6K and 10K examples respectively. We select data samples from the data pool $X_{sota}$. We adopt MT-Bench, AlpacaEval (Li et al., 2023c), and the Open LLM Leaderboard (Beeching et al., 2023) as our benchmarks for automatic evaluation. The Open LLM Leaderboard consists of four classification tasks: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), and TruthfulQA (Lin et al., 2022). We also provide human evaluation results in Appendix D. We first take LLaMA-1-13B as the backbone model and compare Deita with other data selection approaches such as LIMA (Zhou et al., 2023), Alpagasus (Chen et al., 2023), and TAGLM (Lu et al., 2023). Then we compare Deita models based on LLaMA-1-13B, LLaMA-2-13B (Touvron et al., 2023b), and Mistral-7B (Jiang et al., 2023) with other top-performing open-source models, such as Vicuna (Chiang et al., 2023), WizardLM (Xu et al., 2023), Mistral-Instruct (Jiang et al., 2023), and Zephyr (Tunstall et al., 2023). In this paper, we mainly focus on models aligned using SFT without preference training since our contributions lie on data selection in the SFT stage, while we also run direct preference optimization (DPO, Rafailov et al. (2023)) on top of our best SFT model to report stronger performance for reference points. For DPO training, we randomly sample 10K comparison data pairs used in Zephyr that is originally obtained from the UltraFeedback dataset (Cui et al., 2023).
3.3 Results
@@
[…]
@@
We compare Deita to other SOTA models and the random data selection baselines (Random-Select) on Open LLM Leaderboard, results are shown in Table 7. Our Deita SFT models – while trained with only 6K or 10K data – achieve the best average results among SFT-aligned models across different backbone models. Further DPO training greatly boosts the performance of Deita-Mistra-7B by around 5 points on average and helps surpass the Zephyr baseline.
@@
Figure 2: Data-scaling results on MT-Bench. The X-axis represents the # samples used for instruction tuning.
Figure 2: Data-scaling results on MT-Bench. The X-axis represents the # samples used for instruction tuning.

### v2

@@ 1 Introduction
Figure 1: Illustration of the data selection approach. We measure data from three dimensions: complexity, quality, and diversity. $I$ and $R$ represent instruction and response respectively. For Evol Complexity and Evol Quality, we first collect samples with varying complexities or qualities through adopting an evolution-based approach following Xu et al. (2023), then we ask ChatGPT (The term “ChatGPT” in this paper refers to gpt-3.5-turbo-0613) to rank and score the variants of the same data sample for a small seed dataset, and we train our own complexity and quality scorers based on these scores. In the last step, we utilize the trained scorers and adopt a score-first, diversity-aware approach to select the “good” data samples, as we detail in §3.
In the development of large language models (LLMs), aligning LLMs with human preferences is a necessary step for the models to accurately understand human instructions and generate relevant responses. Standard approaches for LLM alignment encompass instruction tuning and reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022). Instruction tuning, or supervised fine-tuning (SFT), refines the pre-trained model using annotated instructional data, often serving as the foundational step before RLHF to facilitate the model’s initial alignment (Touvron et al., 2023b). RLHF, on the other hand, leverages reinforcement learning to train models based on annotated feedback on their generated responses. While RLHF has underpinned the development of ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023), recent researches suggest that instruction tuning in isolation can offer competitive results (Sun et al., 2024; Zhou et al., 2023).
@@
In this work, we seek to define the characteristics of “good data” for instruction tuning, based on which we aim to push data efficiency of instruction tuning further in an automatic manner. To this end, we first explore various methods to quantitatively assess data examples from three key dimensions: complexity, quality, and diversity. Our hypothesis posits that the most effective datasets for instruction tuning are those that are complex, of high quality, and diverse. We examine a wide range of data measurement baselines and propose new metrics, which show a stronger correlation with alignment performance after instruction tuning. Drawing inspiration from Xu et al. (2023), we introduce Evol Complexity and Evol Quality, which evolve a single data point to produce a series of examples varying in complexity or quality, then we rank and score a small subset of them with ChatGPT and train a scorer to predict these scores, as illustrated in Figure 1. Such an evolution-based method automatically yields examples sorted based on complexity or quality, enabling enhanced and finer-grained data measurements. Combined with our diversity metric via distance of model embeddings, we design a simple strategy to select the most effective data examples from a large data pool. Being simple and effective, our data selection strategy paves the way for a paradigm of data-efficient instruction tuning – where fewer training samples can be automatically selected out to yield performance on par with, or even surpassing, models trained on significantly larger datasets.
We present Deita (short for Data-Efficient Instruction Tuning for Alignment), a family of models fine-tuned from the LLaMA (Touvron et al., 2023a; b) and Mistral (Jiang et al., 2023) using our proposed technique to maximize data efficiency. On experiments in terms of MT-bench (Zheng et al., 2023), AlpacaEval (Li et al., 2023c), and the Open LLM Leaderboard (Beeching et al., 2023) demonstrate that Deita is able to outperform or be on par with state-of-the-art instruction following models such as Zephyr (Tunstall et al., 2023), Vicuna (Chiang et al., 2023) and WizardLM (Xu et al., 2023), while using over 10x fewer automatically selected data examples. For example, Deita-Mistral-7B – our model based on Mistral-7B – achieves 7.22 MT-bench and 80.78% AlpacaEval when trained on only 6K data samples with vanilla SFT training. After equipped with direct preference optimization (DPO,  Rafailov et al. (2023)), Deita-Mistral-7B + DPO, trained with 6K SFT and 10K DPO samples, obtains 7.55 MT-Bench and 90.06% AlpacaEval. In addition to the Deita model checkpoints, we release our light yet effective SFT datasets to facilitate alignment for future researches.
@@ 2.2 Experimental Setup
To investigate data selection from large data pools, we construct two data pools with distinct properties to mimic different practical settings: (1) $X_{sota}$, which is constructed by ensembling the training datasets of the state-of-the-art aligned LLMs. This represents the setting where a data pool that is relatively complex, diverse, and of high-quality is available, and we aim to further improve the data efficiency in this case. Specifically, we follow Lu et al. (2023) to combine the datasets WizardLM (Alpaca), WizardLM (ShareGPT), UltraChat (Ding et al., 2023), and ShareGPT (Chiang et al., 2023), resulting in a dataset of  300K samples; and (2) $X_{base}$, which mimics the scenario where the available data pool is overall lower-quality and redundant. This may be closer to the actual setting that people would encounter in practice. For $X_{base}$, we have utilized data from Alpaca (Taori et al., 2023), Dolly (Conover et al., 2023), OAssit (Köpf et al., 2023), and FLAN 2022 (Longpre et al., 2023) to construct a dataset of 100K samples. Such quality-based separation of these open-source instruction tuning datasets roughly aligns with the analysis results in Lu et al. (2023); Li et al. (2023b). Statistics of the two data pools are summarized in Table 1.
@@
In the study of this section, we fine-tune the LLaMA-1 13B on the instruction tuning dataset unless otherwise specified. We assume a data budget $m$ of 6K. We use the same hyperparameters for all as detailed in Appendix A. To evaluate alignment performance, we utilize MT-Bench, a challenging benchmark that is commonly adopted to assess the instruction-following ability. MT-bench consists of multi-turn conversations across various domains such as writing, reasoning, math, and coding. GPT-4 is employed as the judge to score model responses in MT-bench, which was found to produce a high agreement with humans (Zheng et al., 2023).
@@
Table 2: MT-bench of different complexity metrics. All methods select 6K samples. “Pool=50K” denotes the data selection is conducted in a 50K-sized subset due to the cost of using ChatGPT to annotate the entire pool. We include the results of our method on the 50K data pool to make a fair comparison with baselines.
@@ 2.3 From the Complexity Perspective – Evol Complexity
It is often believed that long, difficult, and complex data samples are more beneficial for instruction tuning (Zhao et al., 2023; Cao et al., 2023). For example, Xu et al. (2023) prompt ChatGPT to “evolve” data samples to deliberately increase their complexities, which leads to the state-of-the-art open-source alignment models WizardLM. In this section, we systematically study various metrics to assess the complexity of data and aim to identify the notion of complexity that contributes the most to the instruction-following ability. Concretely, we only consider the complexity dimension and define the selection strategy $\pi_{complexity}$ as selecting the $m$ examples with the highest complexity scores.
@@
As illustrated in the left part of Figure  1, we then ask ChatGPT to rank and score these 6 samples (prompt in Appendix E.2), obtaining the complexity scores $c$ corresponding to the instructions. We emphasize that, distinct from direct scoring, we give ChatGPT all 6 samples within one prompt – these samples represent different evolution stages of the same original sample and such a scoring scheme helps ChatGPT capture the small complexity differences among them, which leads to complexity scores to achieve finer-grained complexity differentiation among samples. We find that this is critical since otherwise, ChatGPT tends to assign similar scores to most examples, as we showcase in Appendix B. After obtaining ChatGPT scores on the small seed dataset, we use the scores to train an LLaMA-1 7B model to predict the complexity score given the input instruction. In multi-turn dialogues, we score each turn separately and use the sum of them as the final score. Across this paper, we use 2K examples randomly sampled from the Alpaca (Taori et al., 2023) as the seed dataset.
@@ Results:
Table 2 presents the results of selecting the 6K data samples from $X_{sota}$ and $X_{base}$ using various complexity metrics. Our Evol-Complexity leads to the best alignment performance. We observe that while Instag Complexity performs well on $X_{sota}$, its performance on $X_{base}$ is only slightly better than Random Selection. In contrast, Evol Complexity achieves superior performance on both datasets, indicating strong robustness across different dataset pools. Our method also has significant advantages compared to methods that heavily rely on annotating with ChatGPT, such as Direct Scoring and Instruction Node. The results also suggest that instruction lengths are not good indicators of preferred data by alignment. Interestingly, perplexity, an intuitive measure for complexity, produces much worse results than the random selection baseline. Through further investigation, we found that samples with large perplexity typically exhibit very short responses.
@@ 2.4 From the Quality Perspective – Evol Quality
In a manner akin to Evol Complexity, we introduce Evol Quality to augment the discernment of quality measurement. For a given data sample $(I_{k}^{(0)},R_{k}^{(0)})$, we prompt ChatGPT to elevate the quality of the response in an evolved way (Appendix E.4 for the prompt). This primarily involves enhancing helpfulness, augmenting relevance, enriching depth, fostering creativity, and supplying additional details. After $M$ iterations, for the same instruction $I_{k}^{(0)}$, we procure a set of responses spanning various qualities for $R_{k}$, denoted as $\{R_{k}^{(0)},\cdots,R_{k}^{(M)}\}$. Similar to Evol Complexity we set $M$ to 5.
@@ Results:
Table 3: MT-bench of different quality measurements. All methods select 6K samples for training. “Pool=50K” denotes the data selection procedure is conducted in a 50K-sized subset due to the cost of ChatGPT to annotate the entire pool. We include the results of our method on the 50K data pool to make a fair comparison with the baselines.
@@ 2.5 From the Diversity Perspective – An Embedding-based Approach
We examine our embedding-based approach, which we refer to as Repr Filter. Specifically, we take the distance between sample $x_{i}$ and its nearest neighbor in $S$ as the metric to define $\mathcal{F}$. We leverage the LLaMA-1 13B model to encode the sentence and compute the cosine distance $d$, then $\mathcal{F}\coloneqq d<\tau$, where $\tau\in(0,1)$ is a threshold hyperparameter. This means that we consider an example $x_{i}$ could increase the diversity of $S$ when the embedding distance between $x_{i}$ and its nearest neighbor is smaller than a threshold. During data selection, we first sort the data pool $X$ according to complexity and quality scores that we will detail in §3.1, then we examine each sample one by one and put $x_{i}$ to $S$ if $d<\tau$, where $S$ is initialized as empty. This process stops until the size of $S$ reaches the data budget $m$. We set threshold $\tau$ as $0.9$ in all the relevant experiments across this paper, while we provide analysis results on $\tau$ and different sentence representations in Appendix C.1.
@@ 3.1 Method
We train Deita using the selected dataset with $m$ samples. We denote the resulting models as $\textsc{Deita}_{m}$. In this paper, we train Deita models based on LLaMA-1-13B, LLaMA-2-13B, and Mistral-7B respectively, and the training details are described in Appendix A.
3.2 Experimental Setup
We train the Deita models with data budget of 6K and 10K examples respectively. We select data samples from the data pool $X_{sota}$. We adopt MT-Bench, AlpacaEval, and the Open LLM Leaderboard as our benchmarks for automatic evaluation. The Open LLM Leaderboard consists of four classification tasks: ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), and TruthfulQA (Lin et al., 2022). We also provide human evaluation results in Appendix D. We first take LLaMA-1-13B as the backbone and compare Deita with other data selection approaches such as LIMA (Zhou et al., 2023), Alpagasus (Chen et al., 2023), and TAGLM (Lu et al., 2023). Then we compare Deita models based on LLaMA-1-13B, LLaMA-2-13B, and Mistral-7B with other top-performing open-source models, such as Vicuna, WizardLM, Mistral-Instruct (Jiang et al., 2023), and Zephyr. In this paper, we mainly focus on models aligned using SFT without preference training since our contributions lie on data selection in the SFT stage, while we also run direct preference optimization (DPO, Rafailov et al. (2023)) on top of our best SFT model to report stronger performance for reference points. For DPO training, we randomly sample 10K comparison data pairs used in Zephyr that is originally obtained from the UltraFeedback dataset (Cui et al., 2023).
3.3 Results
@@
Deita-LLaMA1-13B6K 6K ${\bf 6.46}$ ${\bf 77.08}$
Table 5: Comparison of different data selection approaches, the backbone is LLaMA-1-13B. For Alpagasus, we could not use ChatGPT to score all the examples in the pool due to the cost, thus we score 50K random examples and select. $\dagger$ denotes the results obtained by using their released LLaMA-7B tagger model for a fair comparison.
@@
Deita-Mistral-7B6K 6K / SFT 7.22 80.78
Deita-Mistral-7B10K 10K / SFT 7.32 81.67
Deita-Mistral-7B6K + DPO 6K / SFT +10K / DPO 7.55 90.06
@@
We first compare Deita with other data selection approaches in Table 5, where Deita-LLaMA16K clearly outperforms other approaches by a large margin. In Table 6, we train LLaMA-1-13B, LLaMA-2-13B, and Mistral-7B with our selected 6K and 10K data respectively, and compare with other state-of-the-art SFT models as well as random data selection baselines (Random-Select). Across all three backbone models, the SFT aligned Deita models outperform almost all other SFT-aligned models. Deita models based on LLaMA-2 even outperforms LLaMA2-13B-Chat that undergoes RLHF training with carefully crafted human annotations. Notably, Deita-Mistral-7B10K based on Mistral-7B achieves a 7.32 MT-Bench score, which is the state-of-the-art result among all open-source SFT models at 7B and 13B sizes. In the meanwhile, we note that the gains of Deita on AlpacaEval are not apparently consistent with the gains on MT-Bench. To analyze such difference between MT-Bench and AlpacaEval, we plot the standard radar plot of the 8 substasks on MT-Bench, as shown in Figure 3. It clearly demonstrates that the Deita Mistral models achieve high MT-Bench scores due to the enhanced performance on advanced abilities such as coding, math, and reasoning, which are not prominent in AlpacaEval. When equipped with DPO training, our Deita-Mistral-7B10K+DPO variant achieves 7.55 MT-Bench and 90.06% AlpacaEval scores, which is comparable to zephyr-beta that trains on 30x more data, and slightly lags behind the recent Mistral-7B-Instruct-v0.2 model whose alignment approach and data are not public.
@@
We compare Deita to other SOTA models and the random data selection baselines (Random-Select) on Open LLM Leaderboard, results are shown in Table 7. Our Deita SFT models – while trained with only 6K or 10K data – achieve the best average results among SFT-aligned models across different backbones. Further DPO training greatly boosts the performance of Deita-Mistra-7B by around 5 points on average and helps surpass the Zephyr.
@@
Figure 2: Data-scaling results on MT-Bench. The X-axis represents the # samples used.